please help have admission test for new job in 2hours.
A=B+3(4-C)
if B=5 and C=2
: please help have admission test for new job in 2hours.
A=B+3(4-C)
if B=5 and C=2

Answer by vleith(1174)

A=B+3(4-C)
if B=5 and C=2
A = 5 +3(4-2)
A = 5 +3(2)
A = 5 + 6
A = 11